OpenStudy (anonymous):

integral of (squrt(x) +4)/(x^2) from one to infinity...please i need someone em othain the mothods/process :)) pleaseeeee

OpenStudy (espex):

Would it help you to look at it like this: \[\int\limits_{1}^{\infty}(x+4)^{\frac{1}{2}}(x^{-2})\] Then you could use your power rule and your chain rule on each in conjunction with your product rule.

OpenStudy (anonymous):

oh, thank you verrryyyyyy much :)